OpenStudy (anonymous):

find a rectangular equation for the curve: r= 8sec(θ)/(5csc(θ)-6)

OpenStudy (anonymous):

Try performing the substitutions \(r=\sqrt{x^2+y^2}\) and \(\theta=\arctan\frac{y}{x}\) and simplifying.